# pySLAM: An Open-Source, Modular, and Extensible Framework for SLAM

### v2

@@ Abstract
pySLAM is an open-source Python framework for Visual SLAM, supporting monocular, stereo, and RGB-D cameras. It provides a flexible interface for integrating both classical and modern local features, making it adaptable to various SLAM tasks. The framework includes different loop closure methods, a volumetric reconstruction pipeline, and support for depth prediction models. Additionally, it offers a suite of tools for visual odometry and SLAM applications. Designed for both beginners and experienced researchers, pySLAM encourages community contributions, fostering collaborative development in the field of Visual SLAM.
Introduction
@@
The goal of this document is to present the pySLAM framework, its main features, and usage. pySLAM is a python implementation of a Visual SLAM pipeline that supports monocular, stereo and RGBD cameras. It provides the following features:
@@
  • Integration of depth prediction models within the SLAM pipeline. TThese include DepthPro, DepthAnythingV2, RAFT-Stereo, CREStereo, MASt3R, MVDUSt3R, etc.
  • A collection of other useful tools for VO and SLAM.
A convenient entry-point are the following main scripts:
@@
pySLAM can be used as flexible baseline framework to experiment with VO/SLAM techniques, local features, descriptor aggregators, global descriptors, volumetric integration and depth prediction. It allows to quickly explore, prototype and develop VO/SLAM pipelines. Users should note that pySLAM is a research framework and a work in progress. It is not optimized for real-time performances.
Enjoy it!
System overview
@@
SLAM Workflow
SLAM Components
Note: In some case, I used Processes instead of Threads because in Python 3.8 (used by pySLAM) the Global Interpreter Lock (GIL) allows only one thread can execute at a time within a single process. Multiprocessing avoids this limitation and enables better parallelism, though it involves data duplication via pickling. See this related nice post.
Main System Components
@@
Usage
Once you have run the script install_all_venv.sh / install_all_conda.sh (follow the instructions above according to your OS), you can open a new terminal and start testing the basic Visual Odometry (VO):
. pyenv-activate.sh   #  Activate pyslam python virtual environment. This is only needed once in a new terminal.
./main_vo.py
This will process a default KITTI video (available in the folder data/videos) by using its corresponding camera calibration file (available in the folder settings), and its groundtruth (available in the same data/videos folder). If matplotlib windows are used, you can stop main_vo.py by focusing/clicking on one of them and pressing the key ‘Q’. As explained above, this very basic script main_vo.py strictly requires a ground truth. Now, with RGBD datasets, you can also test the RGBD odometry with the classes VisualOdometryRgbd or VisualOdometryRgbdTensor (ground truth is not required here).
@@
. pyenv-activate.sh   #  Activate pyslam python virtual environment. This is only needed once in a new terminal.
./main_slam.py
This will process the same default KITTI video (available in the folder data/videos) by using its corresponding camera calibration file (available in the folder settings). You can stop it by focusing/clicking on one of the opened windows and pressing the key ‘Q’ or closing the 3D pangolin GUI.
With both scripts, in order to process a different dataset, you need to update the file config.yaml: * Select your dataset type in the section DATASET (further details in the section Datasets below for further details). This identifies a corresponding dataset section (e.g. KITTI_DATASET, TUM_DATASET, etc). * Select the sensor_type (mono, stereo, rgbd) in the chosen dataset section.
* Select the camera settings file in the dataset section (further details in the section Camera Settings below). * The groudtruth_file accordingly (further details in the section Datasets below and check the files io/ground_truth.py and io/convert_groundtruth.py).
Feature tracking
@@
. pyenv-activate.sh   #  Activate pyslam python virtual environment. This is only needed once in a new terminal.
\$./main_feature_matching.py
@@
Some basic examples are available in the subfolder test/loopclosing. In particular, as for feature detection/description, you may want to take a look at test/cv/test_feature_manager.py too.
Loop closing
Different loop closing methods are available, combining aggregation methods and global descriptors.
While running full SLAM, loop closing is enabled by default and can be disabled by setting kUseLoopClosing=False in config_parameters.py. Configuration options can be found in loop_closing/loop_detector_configs.py.
Examples: Start with the examples in test/loopclosing, such as test/loopclosing/test_loop_detector.py.
@@
DBoW2, DBoW3, and VLAD require pre-trained vocabularies. ORB-based vocabularies are automatically downloaded in the data folder (see loop_closing/loop_detector_configs.py).
@@
Most methods do not require pre-trained vocabularies. Specifically: - iBoW and OBindex2: These methods incrementally build bags of binary words and, if needed, convert (front-end) non-binary descriptors into binary ones. - Others: Methods like HDC_DELF, SAD, AlexNet, NetVLAD, CosPlace, and EigenPlaces directly extract global descriptors and process them using dedicated aggregators, independently from the used front-end descriptors.
@@
Double-check your loop detection configuration and verify vocabulary compability
@@
    The back-end and the front-end are using the same descriptor type (this is also automatically checked for consistency).
@@
    A corresponding pre-trained vocubulary is available.
For more details, refer to the vocabulary management section.
If you lack a compatible vocabulary for the selected front-end descriptor type, no loop closing method will be used. You have the following options:
  1. a.
    Create and load the vocabulary (refer to the vocabulary management section).
  2. b.
    Choose an *_INDEPENDENT loop detector method, which works with an independent local_feature_manager.
  3. c.
    Select a vocabular-free loop closing method.
See the file loopclosing/loop_detector_configs.py for further details.
Volumetric reconstruction
@@
The SLAM back-end hosts a volumetric reconstruction pipeline. This is disabled by default. You can enable it by setting kUseVolumetricIntegration=True and selecting your preferred method kVolumetricIntegrationType in config_parameters.py. At present, two methods are available: TSDF and GAUSSIAN_SPLATTING (see dense/volumetric_integrator_factory.py). Note that you need CUDA in order to run GAUSSIAN_SPLATTING method.
At present, the volumetric reconstruction pipeline works with: - RGBD datasets - When a depth estimator is used in the back-end or front-end and a depth prediction/estimation gets available for each processed keyframe.
If you want a mesh as output then set kVolumetricIntegrationExtractMesh=True in config_parameters.py.
@@
Use the script main_map_dense_reconstruction.py to reload a saved sparse map and to perform dense reconstruction by using its posed keyframes as input. You can select your preferred dense reconstruction method directly in the script.
@@
If you are targeting volumetric reconstruction while running SLAM, you can enable a keyframe generation policy designed to manage the spatial distribution of keyframe field-of-view (FOV) centers. The FOV center of a camera is defined as the backprojection of its image center, calculated using the median depth of the frame. With this policy, a new keyframe is generated only if its FOV center is farther than a predefined distance from the nearest existing keyframe’s FOV center. You can enable this policy by setting the following parameters in the yaml setting:
KeyFrame.useFovCentersBasedGeneration: 1  # compute 3D fov centers of camera frames by using median depth and use their distances to control keyframe generation
KeyFrame.maxFovCentersDistance: 0.2       # max distance between fov centers in order to generate a keyframe
@@
The available depth prediction models can be utilized both in the SLAM back-end and front-end. - Back-end: Depth prediction can be enabled in the volumetric reconstruction pipeline by setting the parameter kVolumetricIntegrationUseDepthEstimator=True and selecting your preferred kVolumetricIntegrationDepthEstimatorType in config_parameters.py. - Front-end: Depth prediction can be enabled in the front-end by setting the parameter kUseDepthEstimatorInFrontEnd in config_parameters.py. This feature estimates depth images from input color images to emulate a RGBD camera. Please, note this functionality is still experimental at present time.
Refer to the file depth_estimation/depth_estimator_factory.py for further details. Both stereo and monocular prediction approaches are supported. You can test depth prediction/estimation by using the script main_depth_prediction.py.
Notes: * In the case of a monocular SLAM configuration, do NOT use depth prediction in the back-end volumetric integration: The SLAM (fake) scale will conflict with the absolute metric scale of depth predictions. With monocular datasets, enable depth prediction to run in the front-end. - The depth inference may be very slow (for instance, with DepthPro it takes ~1s per image on my machine). Therefore, the resulting volumetric reconstruction pipeline may be very slow.
Saving and reloading

### v3

@@ Abstract
pySLAM is an open-source Python framework for Visual SLAM that supports monocular, stereo, and RGB-D camera inputs. It offers a flexible and modular interface, integrating a broad range of both classical and learning-based local features. The framework includes multiple loop closure strategies, a volumetric reconstruction pipeline, and support for depth prediction models. It also offers a comprehensive set of tools for experimenting with and evaluating visual odometry and SLAM modules. Designed for both beginners and experienced researchers, pySLAM emphasizes rapid prototyping, extensibility, and reproducibility across diverse datasets. Its modular architecture facilitates the integration of custom components and encourages research that bridges traditional and deep learning-based approaches. Community contributions are welcome, fostering collaborative development and innovation in the field of Visual SLAM. This document [^1] You may find an updated version of this document at:
github.com/luigifreda/pyslam/blob/master/docs/tex/document.pdf
presents the pySLAM framework, outlining its main components, features, and usage.
1 Introduction
pySLAM is an open-source framework designed to accelerate research and development in Visual SLAM (Simultaneous Localization and Mapping) and Spatial AI. It originated from a practical need identified during the development of a SLAM/Spatial AI course: The absence of a unified, accessible, and extensible SLAM ecosystem in C++ and Python. The project was driven by several foundational motivations, including the desire to rapidly prototype and evaluate both classical and learning-based features within complete VO/SLAM pipelines, ensure reproducibility across a wide range of datasets, and significantly lower the entry barrier for newcomers to the field.
A key decision behind pySLAM development was the choice of Python as its primary language. Compared to C++, Python enables faster experimentation and benefits from a mature and expanding ecosystem of deep learning tools and research-oriented APIs. This choice aligns with current trends in computer vision and robotics research, where rapid prototyping and flexible experimentation are essential.
However, designing a modern SLAM framework in Python presents several challenges. The SLAM landscape is highly fragmented – across feature types, datasets, programming languages (Python vs. C++), and development environments (virtual environments, Docker, versioning issues). Deep learning integration further complicates matters due to backend fragmentation (e.g., PyTorch vs. TensorFlow). These inconsistencies often lead to reproducibility bottlenecks and high barriers to experimental research, where setting up a functional baseline can become more laborious than the research itself.
To address these issues, pySLAM is built on several key design principles. Modularity is central: components such as local/global features, loop closure mechanisms, semantic mapping, and volumetric integration can be easily swapped or extended. This architecture supports the use of interchangeable baselines, enabling researchers to plug-and-play with different configurations depending on their goals. Extensibility is facilitated through a base-class architecture that allows new components to be integrated with minimal friction, making pySLAM a flexible platform for benchmarking and algorithm development.
In terms of usability, pySLAM emphasizes simplicity in the build and deployment process.. It supports a unified Python environment compatible with conda, pyenv, and pixi, and offers a one-command installation process that covers setup and dataset acquisition. Cross-platform compatibility (Linux, macOS, Windows with WSL2) and full source accessibility further enhance its utility for both academic and industrial use.
Another cornerstone of the framework is standardized dataset handling. By providing a unified interface to multiple datasets, pySLAM encourages reproducibility and reusability. It supports consistent data loading, map state saving/loading, and automatic generation of evaluation metrics – including final and online trajectory estimation – thus enabling robust comparison between algorithms.
Functionally, pySLAM supports the full visual SLAM pipeline with monocular, stereo, and RGB-D configurations. It accommodates both traditional local features and deep-learning-based methods, and includes loop closure detection integrated with optimization backends like g2o and GTSAM. The framework also supports 3D reconstruction via TSDF and emerging techniques like Gaussian Splatting, and integrates deep learning models for depth prediction and semantic scene understanding.
Together, these capabilities position pySLAM as a powerful research toolkit – aiming to be a “Swiss army knife“ for SLAM development, evaluation, and education. By streamlining reproducibility, supporting modular experimentation, and lowering the technical threshold for entry, pySLAM addresses some of the most persistent challenges in contemporary SLAM research.
The objective of this document is to present the pySLAM framework, its main features, and usage [^2] You may find an updated version of this document at:
github.com/luigifreda/pyslam/blob/master/docs/tex/document.pdf
2 Main Features
@@
pySLAM provides a python implementation of a Visual SLAM pipeline that supports monocular, stereo and RGBD cameras. It offers the following features in a single python environment:
@@
  • Integration of depth prediction models within the SLAM pipeline. These include DepthPro, DepthAnythingV2, RAFT-Stereo, CREStereo, MASt3R, MVDUSt3R, etc.
  • A suite of segmentation models for semantic understanding of the scene, such as DeepLabv3, Segformer, and dense CLIP.
  • Additional tools for VO (Visual Odometry) and SLAM, with built-in support for both g2o and GTSAM, along with custom Python bindings for features not available in the original libraries
  • Built-in support for over 10 dataset types. The list is currently growing.
pySLAM serves as a flexible baseline framework to experiment with VO/SLAM techniques, local features, descriptor aggregators, global descriptors, volumetric integration, depth prediction and semantic mapping. It allows to explore, prototype and develop VO/SLAM pipelines. pySLAM is a research framework and a work in progress. It is not optimized for real-time performance.
3 Overview
3.1 Main Scritps
A convenient entry point is the following collection of main scripts:
@@ 3.1 Main Scritps
  • main_slam_evaluation.py enables automated SLAM evaluation by executing main_slam.py across a collection of datasets and configuration presets (see here).
Figure 1: SLAM workflow.
Figure 2: SLAM components.
3.2 System overview
@@
SLAM Workflow and Components
Fig. 1 illustrates the SLAM workflow, which is composed of six main parallel processing modules:
  • Tracking: estimates the camera pose for each incoming frame by extracting and matching local features to the local map, followed by minimizing the reprojection error through motion-only Bundle Adjustment (BA). It includes components such as pose prediction (or relocalization), feature tracking, local map tracking, and keyframe decision-making.
  • Local Mapping: updates and refines the local map by processing new keyframes. This involves culling redundant map points, creating new points via temporal triangulation, fusing nearby map points, performing Local BA, and pruning redundant local keyframes.
  • Loop Closing: detects and validates loop closures to correct drift accumulated over time. Upon loop detection, it performs loop group consistency checks and geometric verification, applies corrections, and then launches Pose Graph Optimization (PGO) followed by a full Global Bundle Adjustment (GBA). Loop detection itself is delegated to a parallel process, the Loop Detector, which operates independently for better responsiveness and concurrency.
  • Global Bundle Adjustment: triggered by the Loop Closing module after PGO, this step globally optimizes the trajectory and the sparse structure of the map to ensure consistency across the entire sequence.
  • Volumetric Integration: uses the keyframes, with their estimated poses and back-projected point clouds, to reconstruct a dense 3D map of the environment. This module optionally integrates predicted depth maps and maintains a volumetric representation such as a TSDF [15] or incremental Gaussian Splatting-based volume [37, 21].
  • Semantic Mapping: enriches the SLAM map with dense semantic information by applying pixel-wise segmentation to selected keyframes. Semantic predictions are fused across views to assign semantic labels or descriptors to keyframes and map points. The module operates in parallel, consuming keyframes and associated image data from a queue, applying a configured semantic segmentation model, and updating the map with fused semantic features. This enables advanced downstream tasks such as semantic navigation, scene understanding, and category-level mapping.
The first four modules follow the established PTAM [22] and ORB-SLAM [43] paradigm. Here, the Tracking module serves as the front-end, while the remaining modules operate as part of the back-end.
In parallel, the system constructs two types of maps:
  • a sparse map ${\cal M}_{s}=({\cal K},{\cal P})$, composed of a set of keyframes $\cal K$ and 3D points ${\cal P}$ derived from matched features;
  • a volumetric map (or dense map) ${\cal M}_{v}$, constructed by the Volumetric Integration module, which fuses back-projected point clouds from the keyframes $\cal K$ into a dense 3D model.
To ensure consistency between the sparse and volumetric representations, the volumetric map is updated or re-integrated whenever global pose adjustments occur (e.g., after loop closures).
Fig. 2 details the internal components and interactions of the above modules. In certain cases, processes are employed instead of threads. This is due to Python’s Global Interpreter Lock (GIL), which prevents concurrent execution of multiple threads in a single process. The use of multiprocessing circumvents this limitation, enabling true parallelism at the cost of some inter-process communication overhead (e.g., via pickling). For an insightful discussion, see this related post.
3.3 Main System Components
Figure 3: Feature tracker.
@@ 3.3 Main System Components
The Feature Tracker consists of the following key sub-components:
  • Feature Manager: Manages local feature operations. It includes the FeatureDetector, FeatureDescriptor, and adaptors for pyramid management and image tiling.
    • Feature Detector: Identifies salient and repeatable keypoints in the image, such as corners or blobs, which are likely to be robust under viewpoint and illumination changes.
    • Feature Descriptor: Computes a distinctive descriptor for each detected keypoint, encoding its local appearance to enable robust matching across frames. Examples include ORB [54], SIFT [31], or SuperPoint [14] descriptors.
  • Feature Matcher: Establishes correspondences between features in successive frames (or stereo pairs) by comparing their descriptors or directly inferring matches from image content. Matching can be performed using brute-force, k-NN with ratio test, or learned matching strategies. Refere to Sect. 3.3 for futher details.
Sect. 5.1 lists the supported local feature extractors and detectors.
The diagram in Fig. 3 presents the architecture of the Feature Tracker system. It is structured around a feature_tracker_factory, which instantiates specific tracker types such as LK, DES_BF, DES_FLANN, XFEAT, LIGHTGLUE, and LOFTR. Each tracker type creates a corresponding implementation (e.g., LKFeatureTracker, DescriptorFeatureTracker, etc.), all of which inherit from a common FeatureTracker interface.
The FeatureTracker class is composed of several key sub-components, including a FeatureManager, FeatureDetector, FeatureDescriptor, PyramidAdaptor, BlockAdaptor, and FeatureMatcher. The FeatureManager also encapsulates instances of the detector, descriptor, and adaptors, highlighting the modular and reusable design of the tracking pipeline.
Figure 4: Feature matcher.
Figure 5: Loop detector.
@@
The diagram in Fig. 4 illustrates the architecture of the Feature Matcher module. At its core is the feature_matcher_factory, which instantiates matchers based on a specified matcher_type, such as BF, FLANN, XFEAT, LIGHTGLUE, and LOFTR. Each of these creates a corresponding matcher implementation (e.g., BfFeatureMatcher, FlannFeatureMatcher, etc.), all inheriting from a common FeatureMatcher interface.
The FeatureMatcher class encapsulates several configuration parameters and components, including the matcher engine (cv2.BFMatcher, FlannBasedMatcher, xfeat.XFeat, etc.), as well as the matcher_type, detector_type, descriptor_type, norm_type, and ratio_test fields. This modular structure supports extensibility and facilitates switching between traditional and learning-based feature matching backends.
The Section 5.2 reports a list of supported feature matchers.
Figure 6: Depth estimator.
@@
The diagram in Fig. 5 shows the architecture of the Loop Detector component. A central loop_detector_factory instantiates loop detectors based on the selected global_descriptor_type, which may include traditional descriptors (e.g., DBOW2, VLAD, IBOW) or deep learning-based embeddings (e.g., NetVLAD, CosPlace, EigenPlaces).
Each descriptor type creates a corresponding loop detector implementation (e.g., LoopDetectorDBoW2, LoopDetectorNetVLAD), all of which inherit from a base class hierarchy. Traditional methods inherit directly from LoopDetectorBase, while deep learning-based approaches inherit from LoopDetectorVprBase, which itself extends LoopDetectorBase. This design supports modular integration of diverse place recognition techniques within a unified loop closure framework.
The Section 5.3 reports a list of supported loop closure methods with the adopted global descriptors and local descriptor aggregation methods.
@@
The diagram in Fig. 6 illustrates the architecture of the Depth Estimator module. A central depth_estimator_factory creates instances of various depth estimation backends based on the selected depth_estimator_type, including both traditional and learning-based methods such as DEPTH_SGBM, DEPTH_RAFT_STEREO, DEPTH_ANYTHING_V2, DEPTH_MAST3R, and DEPTH_MVDUST3R.
Each estimator type instantiates a corresponding implementation (e.g., DepthEstimatorSgbm, DepthEstimatorCrestereoMegengine, etc.), all inheriting from a common DepthEstimator interface. This base class encapsulates shared dependencies such as the camera, device, and model components, allowing for modular integration of heterogeneous depth estimation techniques across stereo, monocular, and multi-view pipelines.
The Section 5.4 reports a list of supported depth estimation/prediction models.
Figure 7: Volumetric integrator.
@@
The diagram in Fig. 7 illustrates the structure of the Volumetric Integrator module. At its core, the volumetric_integrator_factory generates specific volumetric integrator instances based on the selected volumetric_integrator_type, such as TSDF and GAUSSIAN_SPLATTING.
Each type instantiates a dedicated implementation (e.g., VolumetricIntegratorTSDF, VolumetricIntegratorGaussianSplatting), which inherits from a common VolumetricIntegratorBase. This base class encapsulates key components including the camera, a keyframe_queue, and the volume, enabling flexible integration of various 3D reconstruction methods within a unified pipeline.
The Section 5.5 reports a list of supported volume integration methods.
Figure 8: Semantic Mapping.
Semantic Mapping
The diagram in Fig. 8 outlines the architecture of the Semantic Mapping module. At its core is the semantic_mapping_factory, which creates semantic mapping instances according to the selected semantic_mapping_type. Currently, the supported type is DENSE, which instantiates the SemanticMappingDense class. This class extends SemanticMappingBase and runs asynchronously in a dedicated thread to process keyframes as they become available.
SemanticMappingDense integrates semantic information into the SLAM map by leveraging per-frame predictions from a semantic segmentation model. The segmentation model is instantiated via the semantic_segmentation_factory, based on the selected semantic_segmentation_type. Supported segmentation backends include DEEPLABV3, SEGFORMER, and CLIP, each of which corresponds to a dedicated class (SemanticSegmentationDeepLabV3, SemanticSegmentationSegformer, SemanticSegmentationCLIP) inheriting from the shared SemanticSegmentationBase.
The system supports multiple semantic feature representations – such as categorical labels, probability vectors, and high-dimensional feature embeddings – and fuses them into the map using configurable methods like count-based fusion, Bayesian fusion, or feature averaging.
This modular design decouples semantic segmentation from mapping logic, enabling flexible combinations of segmentation models, datasets (e.g., NYU40, Cityscapes), and fusion strategies. It also supports customization via configuration files or programmatic APIs for dataset-specific tuning or deployment.
The Section 4.3 provides a list of supported semantic segmentation methods. A more in-depth presentation of the semantic mapping module is provided in [41].
4 Usage
Open a new terminal and start experimenting with the scripts. In each new terminal you are supposed to start with this command:
[…]
commandchars=
[…]
The file config.yaml can be used as a unique entry-point to configure the system and its global configuration parameters contained in pyslam/config_parameters.py. Further information on how to configure pySLAM are provided here.
Visual odometry
The basic Visual Odometry (VO) can be run with the following commands:
[…]
commandchars=
[…]
By default, the script processes a KITTI video (available in the folder data/videos) by using its corresponding camera calibration file (available in the folder settings), and its groundtruth (available in the same data/videos folder). If matplotlib windows are used, you can stop main_vo.py by clicking on one of them and pressing the key ‘Q’. As explained above, this very basic script main_vo.py strictly requires a ground truth. Now, with RGBD datasets, you can also test the RGBD odometry with the classes VisualOdometryRgbd or VisualOdometryRgbdTensor (ground truth is not required here).
Full SLAM
@@ 4 Usage
[…]
commandchars=
[…]
This will process the same default KITTI video (available in the folder data/videos) by using its corresponding camera calibration file (available in the folder settings). You can stop it by clicking on one of the opened windows and pressing the key ‘Q’ or closing the 3D pangolin GUI.
Selecting a dataset and different configuration parameters
The file config.yaml can be used as a unique entry-point to configure the system, the target dataset and its global configuration parameters set in pyslam/config_parameters.py. To process a different dataset with both VO and SLAM scripts, you need to update the file config.yaml:
  • Select your dataset type in the section DATASET (further details in the section Datasets below for further details). This identifies a corresponding dataset section (e.g. KITTI_DATASET, TUM_DATASET, etc).
  • Select the sensor_type (mono, stereo, rgbd) in the chosen dataset section.
  • Select the camera settings file in the dataset section (further details in the section Camera Settings below).
  • Set the groudtruth_file accordingly. Further details in the section Datasets below (see also the files pyslam/io/ground_truth.py and pyslam/io/convert_groundtruth_to_simple.py).
You can use the section GLOBAL_PARAMETERS of the file config.yaml to override the global configuration parameters set in pyslam/config_parameters.py. This is particularly useful when running a SLAM evaluation.
4.1 Feature tracking
@@
[…]
commandchars=
[…]
@@
Some basic examples are available in the subfolder test/cv. In particular, as for feature detection/description, you may want to take a look at test/cv/test_feature_manager.py too.
4.2 Loop closing
Many loop closing methods are available, combining different aggregation methods and global descriptors.
While running full SLAM, loop closing is enabled by default and can be disabled by setting kUseLoopClosing=False in pyslam/config_parameters.py. Different configuration options LoopDetectorConfigs can be found in pyslam/loop_closing/loop_detector_configs.py: Code comments provide additional useful details.
One can start experimenting with loop closing methods by using the examples in test/loopclosing. The example test/loopclosing/test_loop_detector.py is the recommended entry point.
@@ 4.2 Loop closing
DBoW2, DBoW3, and VLAD require pre-trained vocabularies. ORB-based vocabularies are automatically downloaded into the data folder (see pyslam/loop_closing/loop_detector_configs.py).
@@
Once you have trained the vocabulary, you can add it in pyslam/loop_closing/loop_detector_vocabulary.py and correspondingly create a new loop detector configuration in pyslam/loop_closing/loop_detector_configs.py that uses it.
@@
Most methods do not require pre-trained vocabularies. Specifically:
  • iBoW and OBindex2: These methods incrementally build bags of binary words and, if needed, convert (front-end) non-binary descriptors into binary ones.
  • Others: Methods like HDC_DELF, SAD, AlexNet, NetVLAD, CosPlace, and EigenPlaces directly extract their specific global descriptors and process them using dedicated aggregators, independently from the used front-end descriptors.
@@
Verify your loop detection configuration and verify vocabulary compability
Loop detection method based on a pre-trained vocabulary
@@
    The back-end and the front-end are using the same descriptor type (this is also automatically checked for consistency) or their descriptor managers are independent (see further details in the configuration options LoopDetectorConfigs available in pyslam/loop_closing/loop_detector_configs.py).
@@
    A corresponding pre-trained vocubulary is available. For more details, refer to the vocabulary management section.
Missing vocabulary for the selected front-end descriptor type
If you lack a compatible vocabulary for the selected front-end descriptor type, you can follow one of these options:
1. Create and load the vocabulary (refer to the vocabulary management section).
2. Choose an *_INDEPENDENT loop detector method, which works with an independent local_feature_manager.
3. Select a vocabulary-free loop closing method.
See the file pyslam/loop_closing/loop_detector_configs.py for further details.
4.3 Volumetric reconstruction
@@
The SLAM back-end hosts a volumetric reconstruction pipeline. This is disabled by default. You can enable it by setting kUseVolumetricIntegration=True and selecting your preferred method kVolumetricIntegrationType in pyslam/config_parameters.py. At present, two methods are available: TSDF and GAUSSIAN_SPLATTING (see pyslam/dense/volumetric_integrator_factory.py). Note that you need CUDA in order to run GAUSSIAN_SPLATTING method.
At present, the volumetric reconstruction pipeline works with:
  • RGBD datasets
  • When a depth estimator is used
    • $\circ$
      in the back-end with STEREO datasets (you can’t use depth prediction in the back-end with MONOCULAR datasets, further details here)
    • $\circ$
      in the front-end (to emulate an RGBD sensor) and a depth prediction/estimation gets available for each processed keyframe.
To obtain a mesh as output, set kVolumetricIntegrationExtractMesh=True in config_parameters.py.
@@
Use the script main_map_dense_reconstruction.py to reload a saved sparse map and perform dense reconstruction by using its posed keyframes as input. You can select your preferred dense reconstruction method directly in the script.
@@
    The directory <gs_checkpoint_path> is expected to have the following structure:
[…]
    commandchars=
[…]
@@
If you are targeting volumetric reconstruction while running SLAM, you can enable a keyframe generation policy designed to manage the spatial distribution of keyframe field-of-view (FOV) centers. The FOV center of a camera is defined as the backprojection of its image center, computed using the median depth of the frame. With this policy, a new keyframe is generated only if its FOV center lies beyond a predefined distance from the nearest existing keyframe’s FOV center. You can enable this policy by setting the following parameters in the yaml setting:
[…]
commandchars=
[…]
@@
The available depth prediction models can be utilized both in the SLAM back-end and front-end.
  • Back-end: Depth prediction can be enabled in the volumetric reconstruction pipeline by setting the parameter kVolumetricIntegrationUseDepthEstimator=True and selecting your preferred kVolumetricIntegrationDepthEstimatorType in pyslam/config_parameters.py.
  • Front-end: Depth prediction can be enabled in the front-end by setting the parameter kUseDepthEstimatorInFrontEnd in pyslam/config_parameters.py. This feature estimates depth images from input color images to emulate a RGBD camera. Please, note this functionality is still experimental at present time [WIP].
Notes:
  • In the case of a monocular SLAM, do NOT use depth prediction in the back-end volumetric integration: The SLAM (fake) scale will conflict with the absolute metric scale of depth predictions. With monocular datasets, you can enable depth prediction to run in the front-end (to emulate an RGBD sensor).
  • Depth inference may be very slow (for instance, with DepthPro it takes ~1s per image on a typical machine). Therefore, the resulting volumetric reconstruction pipeline may be very slow.
Refer to the file pyslam/depth_estimation/depth_estimator_factory.py for further details. Both stereo and monocular prediction approaches are supported. You can test depth prediction/estimation by using the script main_depth_prediction.py.
Semantic mapping
The semantic mapping pipeline can be enabled by setting the parameter kDoSemanticMapping=True in config_parameters.py. The best way of configuring the semantic mapping module used is to modify it in semantic_mapping_configs.py.
Different semantic mapping methods are available Currently, we support semantic mapping using dense semantic segmenetation.
  • DEEPLABV3: from torchvision, pre-trained on COCO/VOC.
  • SEGFORMER: from transformers, pre-trained on Cityscapes or ADE20k.
  • CLIP: from f3rm package for open-vocabulary support.
Semantic features are assigned to keypoints on the image and fused into map points. The semantic features can be:
  • Labels: categorical labels as numbers.
  • Probability vectors: probability vectors for each class.
  • Feature vectors: feature vectors obtained from an encoder. This is generally used for open vocabulary mapping.
4.4 Saving and reloading